OpenStudy (anonymous):

Consider the given equation ln(x^2 y)+2y^6−x^2=7+e^6 , where x > 0 . Assume that it " determines" an implicit differentiation function f such that y= f(x). a) find dy/dx, also known as Y. b) verify that point P(e^3, e) lies on the graph of the given equation.

OpenStudy (anonymous):

\[equation: \ln (x^2 y) + 2 y^6 - x^2 = 7 + e^6\]